Meile Machine​ Shop, Inc., has a​ 1-year contract for the production of

250,000

gear housings for a new​ off-road vehicle. Owner Larry Meile hopes the contract will be extended and the volume increased next year. Meile has developed costs for three alternatives. They are​ general-purpose equipment​ (GPE), flexible manufacturing system (FMS)

LOADING...

​, and​ expensive, but​ efficient, dedicated machine​ (DM). The cost data​ follow:

​General-Purpose Equipment​ (GPE)

Flexible Manufacturing System​ (FMS)

Dedicated Machine​ (DM)

Annual contracted units

250,000         

250,000              

250,000      

Annual fixed cost

$150,000         

$225,000              

$525,000      

Per unit variable cost

$16.00         

$14.75              

$13.00      

Based on the​ cost, the process that is best suited for the current contracted volume is

DM

FMS

GPE

.
Answer #1

INDIFFERENCE POINT = (FIXED COST A - FIXED COST B) / (VARIABLE COST B - VARIABLE COST A)


FC A = 150000
VC A = 16
FC B = 225000
VC B = 14.75
FC C = 525000
VC C = 13


INDIFFERENCE POINT FROM A TO B RANGE = (150000 - 225000) / (14.75 - 16) = 60000

INDIFFERENCE POINT FROM B TO C RANGE = (225000 - 525000) / (13 - 14.75) = 171429

WE WILL CHOOSE DM FOR 250000 UNITS BECAUSE IT HAS A LOWER VARIABLE COST PER UNIT.